A plane leaves Chicago for LAX 2100km away flying at 500 km/hr. One hour later a jet leaves LAX for Chicago flying 200 km/hr faster. If the plane left at noon, at what time do they cross each other?
Answer by JulietG(1812)

C = 500
L = C+ 200
Let x be the time
What you're trying to find is:
2100 = C + x(C+L) [the first C is for the hour head start]
We know the value of L, so we'll substitute it
2100 = C + x(C+C+200)
We know the value of C, so we'll substitute it
2100 = 500 + x(500+500+200)
Add
2100 = 500 +1200x
Subtract 500 from each side
1600 = 1200x
Divide each side by 1200
1.333333 = x
.
The plane leaves Chicago at 12:00 noon. It has an hour head start on the jet. It travels 500 miles before the jet leaves.
At 1:00, the jet leaves LAX. We solved for x earlier, which was the additional time. 1 1/3 hour after 1 pm is 2:20 PM.
At 2:00, jet = 700, plane = 500 + 500
1/3 of an additional hour is 233.33333 for the jet and 166.666667 for the plane. 400 on top of the previously traveled 1700.
Success!